OpenStudy (anonymous):

A baseball team plays in a stadium that holds 66000 spectators. With the ticket price at $11 the average attendance has been 30000. When the price dropped to $9, the average attendance rose to 33000. a) Find the demand function p(x), where x is the number of the spectators. (Assume p(x) is linear.)

OpenStudy (anonymous):

I've found where you take (33000-3000) over (9-11) (it's -1500) and I plugged that back in to 11=-1500(30,000)+b and got b=45000011 but after that I have no idea what to do

OpenStudy (anonymous):

If x is spectators, then your slope should have been delta y / delta x = (11-2)/(30,000-33,000)

OpenStudy (anonymous):

oops 11-9 =2 der

OpenStudy (anonymous):

m=-2/3000=-1/1500

OpenStudy (anonymous):

oh! that makes much more sense, thanks! Silly me. The next part says, "How should ticket prices be set to maximize revenue" Could you help me with that?

OpenStudy (anonymous):

with the new slope, your set-up with the y=mx+b was good 11=(-1/1500)(30000)+b

OpenStudy (anonymous):

thus giving you b=31! got it!

OpenStudy (anonymous):

perfect... now the next part... hmmm...

OpenStudy (anonymous):

is this an elasticity problem?

OpenStudy (anonymous):

here, I'll screenshot the whole thing for you!

OpenStudy (anonymous):

WAIT!

OpenStudy (anonymous):

My professor just emailed us saying the next part won't actually be a correct answer so we need not do it, but thank you so much anyways!

OpenStudy (anonymous):

bummer... I just got it :)

OpenStudy (anonymous):

Well, if you're doing them for fun now, feel free to answer the next question I post then :)

OpenStudy (anonymous):

lol...k... this is business calc. , right?

OpenStudy (anonymous):

Totally just saw this, so sorry! but yes!